Find the two repeated limits of the function f(x ,y) = (y-x/y+x) (1+x^2)/1+y^2) at (0,0) . Does
the simultaneous limit of f exist as (x, y) →(0,0) ? Give reasons for your answer.
1
Expert's answer
2020-03-31T13:32:31-0400

"\lim_{y\rightarrow0} \lim_{x\rightarrow0} \dfrac {y-x} {y+x}\cdot \dfrac {1+x^2}{1+y^2}=
 \lim_{y\rightarrow0} 1\cdot \dfrac {1}{1+y^2}=1"

"\lim_{x\rightarrow0} \lim_{y\rightarrow0} \dfrac {y-x} {y+x}\cdot \dfrac {1+x^2}{1+y^2}=
 \lim_{x\rightarrow0} -1\cdot \dfrac {1+x^2}{1}=-1"

The repeated limits don`t coincide so the simultaneous limit doesn`t exist